# A Closer Look at Model Collapse: From a Generalization-to-Memorization Perspective

### v2

@@ 4 Mitigating Model Collapse with Data Selection Methods
$\max_{{\mathcal{D}}\subset\mathcal{S},\;|{\mathcal{D}}|=N}\hat{H}_{1}(\mathcal{D})\quad\Longleftrightarrow\quad\max_{{\mathcal{D}}\subset\mathcal{S},\;|{\mathcal{D}}|=N}\sum_{\bm{x}\in{\mathcal{D}}}\log\min_{\begin{subarray}{c}\bm{y}\in{\mathcal{D}}\backslash\bm{x}\end{subarray}}\kappa(\bm{x},\bm{y}),$
@@
Threshold Decay Filter is a soft variant of Greedy Selection that provides adjustable control over the selection strength. When the initial threshold is set sufficiently high and the decay factor is close to $1$, the Threshold Decay Filter behaves similarly to Greedy Selection. Conversely, if both the initial threshold and decay factor are set to 0, the filter does not filter out any data point and reduces to the vanilla replace or accumulate-subsample paradigm. In practice, we first extract image features using a DINOv2 (Oquab et al., 2024) model and compute distances in the feature space, i.e., $\kappa(\bm{x},\bm{y})=\|h(\bm{x})-h(\bm{y})\|_{2}$, where $h(\cdot)$ denotes the feature extractor.
@@ 5 Experiments
We empirically evaluate how our data selection strategies introduced in Section˜4 interact with the two self-consuming paradigms: replace and accumulate-subsample. In both settings, our method serves as a plug-in component for selecting high-quality and diverse training samples from the candidate pool. We demonstrate that the proposed strategies effectively alleviate memorization and reduce FID scores, thereby mitigating model collapse. Additionally, experiments on classifier-free guidance (CFG) (Ho and Salimans, 2022) generation show that our method effectively mitigates diversity collapse of CFG.
@@
We conduct experiments on three widely used image generation benchmarks. CIFAR-10 (Krizhevsky et al., 2009) consists of $32\times 32$ color images in $10$ classes. Due to computational constraints, we use a subset of 32,76832{,}768 training images. Our goal is not to achieve state-of-the-art FID among large diffusion models but rather to demonstrate that our method mitigates memorization in the self-consuming loop. As shown in Section˜3, this subset is sufficient to observe the transition from generalization to memorization. We also conduct experiments on subsets of FFHQ (Jain et al., 2004), downsampled to $32\times 32$ resolution, and MNIST (LeCun et al., 1998), using 8,1928{,}192 and 12,00012{,}000 training images, respectively.
@@
For CIFAR-10 and FFHQ, we employ a UNet-based backbone (Ronneberger et al., 2015) designed for $32\times 32$ RGB images, which predicts noise residuals. The network contains approximately 16M parameters. For MNIST, we use a similar UNet-based architecture adapted for single-channel inputs, with a total of 19M parameters. Detailed network configurations are provided in the Appendix.
@@
For iterative training and sampling, our implementation is based on the Hugging Face Diffusers codebase (von Platen et al., 2022) of DDPM. We use a mixed-precision training of FP16 to train the models. We adopt an Adam optimizer with a learning rate of $10^{-4}$ and a weight decay of $10^{-6}$. The batch size is $128$. A $1000$-step denoising process is used, with all other hyperparameters set to their default values. For Threshold Decay Filter, we use an initial threshold of $60$ and a decay rate of $0.95$. We show in the Appendix that our method is robust in a wide range of hyperparameters.
@@
We use the generalization score and entropy to evaluate the effectiveness of our method in mitigating memorization. Additionally, we adopt the Fréchet Inception Distance (FID) (Heusel et al., 2017) as a metric to quantify the distributional divergence between generated images and real images.
@@ Results.
Diversity Improvement on Classifier-Free Diffusion Guidance (CFG) (Ho and Salimans, 2022).
We further validate the effectiveness of our methods in the CFG setting and show that they substantially improve the diversity of generated images. CFG is a widely used conditional generation technique that consistently enhances perceptual quality but often sacrifices diversity (Chidambaram et al., 2024; Ho and Salimans, 2022; Li et al., 2025a). Prior work by Yoon et al. (2024) identifies the CFG scale as a key factor influencing the rate of model collapse and suggests that setting a moderate scale can help mitigate model collapse. Experimentally, we observe that the CFG generation can indeed generate clearer images than the unconditional baseline, as compared in Figures˜9(a) and 9(b). However, the diversity of generated images collapses rapidly even with a modest guidance scale, with samples within each class soon becoming nearly identical. In contrast, Figures˜9(c) and 9(d) demonstrate that augmenting CFG with our data selection methods significantly improves image diversity and yields substantially lower FID scores compared to the vanilla CFG paradigm. These results demonstrate that our approach effectively mitigates the diversity loss of CFG while preserving its quality advantage throughout the self-consuming loop.
@@
Figure 9: Comparisons of unconditional, CFG, and CFG augmented with filter method. (a)-(c): Generated samples on MNIST (LeCun et al., 1998) at the $8$-th iteration under the accumulate paradigm. The FIDs of images in (a)-(c) are $74.4$, $66.2$, and $\bm{22.4}$ respectively. (d): The estimated entropy of the generated dataset over iterations, which reflects the diversity of the generated images.
@@ 6 Conclusion and Discussion
Based on this work, we believe many future directions could be further investigated. This paper doesn’t rigorously explain why the entropy is collapsing. The finite dataset size, training and sampling errors, and the model bias could all contribute to the collapsing entropy (or the information loss). We envision that a more formal theoretical analysis of collapse dynamics could be developed based on theoretical models such as the mixture of Gaussians (Li et al., 2025b; Wang et al., 2024a; Chen et al., 2024). Besides, this framework could be extended to the language modality, investigating the discrete diffusion model (Lou et al., 2024; Rojas et al., 2025; Kim et al., 2025). Additionally, a more efficient algorithm is needed, as the current greedy selection method is computationally expensive.
Acnkowledgement
LS, MW, HZ, ZZ, and QQ acknowledge funding support from NSF CCF-2212066, NSF CCF-2212326, NSF IIS 2402950, and ONR N000142512339, and the Google Research Scholar Award. MT is grateful for partial supports by NSF Grants DMS-1847802, DMS-2513699, DOE Grant DE-NA0004261, and Richard Duke Fellowship. We also thank Prof. Peng Wang (University of Michigan and University of Macau), Mr. Xiang Li (University of Michigan), and Mr. Xiao Li (University of Michigan) for fruitful discussions and valuable feedback.
@@ A.1 Model Collapse
As state-of-the-art generative models continue to improve in image quality, AI-generated images have become increasingly indistinguishable from real ones and are inevitably incorporated into the training datasets of next-generation models. In fact, Alemohammad et al. [2024a] have shown that the LAION-5B dataset [Schuhmann et al., 2022], which is used to train Stable Diffusion, indeed contains synthetic data. Unfortunately, recent studies [Alemohammad et al., 2024b, Dohmatob et al., 2024b, Kazdan et al., 2024, Dohmatob et al., 2024d, a, Alemohammad et al., 2024a, Feng et al., 2024] demonstrate that model performance deteriorates under such recursive training, potentially leading to model collapse, where the model generates increasingly homogeneous and meaningless content. The concept of model collapse was first introduced in Shumailov et al. [2024], which also provides a theoretical framework based on Gaussian models. Their analysis shows that if a Gaussian model is recursively estimated using data generated by its predecessor, its variance converges to zero—collapsing the distribution into a delta function concentrated at a single point.
@@
Following this important line of work, substantial research has further explored the model collapse phenomenon. Suresh et al. [2024] extend the theoretical analysis from Gaussian to Bernoulli and Poisson models. Dohmatob et al. [2024a] study recursive training under high-dimensional linear and ridge regression settings, providing a linear error rate and proposing optimal regularization to mitigate collapse. Dohmatob et al. [2024d] argue that the conventional scaling laws for foundation models break down when synthetic data is incorporated into the training set. Wyllie et al. [2024] demonstrate that model biases are amplified through recursive generation and proposes algorithmic reparation techniques to eliminate such biases and negative semantic shifts. Seddik et al. [2024] introduce a statistical model that characterizes the collapse process in language models and theoretically estimates the maximum allowable proportion of synthetic data before collapse occurs, validated empirically. Ferbach et al. [2024] show that human preferences can be amplified in an iterative loop: modeling human curation as a reward process, the curated distribution $p_{t}$ converges to $p^{*}$ that maximizes the expected reward as $t\to\infty$. In the work of Ferbach et al. [2024], the reward $r(x)$ is a pointwise function over individual images, whereas our method can be viewed as a curation strategy that maximizes dataset-level entropy at each iteration. Whether the theory in Ferbach et al. [2024] extends to dataset-wise rewards $r(S)$, such as dataset entropy, remains an open question. Nonetheless, their intuition aligns with our results that entropy can be iteratively enhanced compared to vanilla methods. Fu et al. [2025] provide the first generalization error bounds for Self-Consuming Training Loops (STLs), showing that mitigating model collapse requires maintaining a non-negligible portion of real data and ensuring recursive stability. In contrast, our work focuses on a specific collapse phenomenon—the generalization-to-memorization transition—and introduces an entropy-based data selection algorithm to mitigate this behavior. Zhu et al. [2024a] provide a rigorous theoretical analysis for the model collapse in rectified flow models and then propose methods to mitigate it. Finally, Cui et al. [2025] investigate collapse in diffusion models by analyzing the training dynamics of a two-layer autoencoder optimized via stochastic gradient descent, showing how network architecture shapes the model collapse.
@@
To mitigate model collapse, several strategies have been proposed in prior work. One common approach is to accumulate all previously generated samples along with real data into the training set—referred to as the accumulate paradigm in our paper. This strategy has been both empirically and theoretically validated. For example, Gerstgrasser et al. [2024] show that the accumulate paradigm prevents divergence of test error under a linear regression setup. Dey and Donoho [2024] further establish the universality of the error upper bound across a broad class of canonical statistical models. In a related setting, Kazdan et al. [2024] study an accumulate-subsample variant, confirms that the test error plateaus and examines interactions between real and synthetic data. Bertrand et al. [2024] provide theoretical guarantees for stability in iterative training, assuming the initial model trained on real data is sufficiently accurate and the clean data proportion in each iteration remains high. Despite these findings, recent work [Dohmatob et al., 2024c] present a robust negative result, showing that model collapse persists even when real and synthetic data are mixed. Similarly, Alemohammad et al. [2024a] argue that the accumulate paradigm merely delays, rather than prevents, collapse. Introducing fresh real data in each iteration may be necessary for long-term robustness. Other complementary approaches include verification [Feng et al., 2024] and re-editing [Zhu et al., 2024b]. For instance, Feng et al. [2024] theoretically underscore the importance of data selection, though it does not propose a specific method. Zhu et al. [2024b] target language models and introduces a token-level editing mechanism with theoretical guarantees. Compared to prior work, this paper proposes a novel entropy-based data selection method for diffusion models that improves both generalizability and image quality, thereby mitigating model collapse.
@@
Over the past few years, extensive research has explored model collapse from various perspectives and dimensions. An important study [Schaeffer et al., 2025] makes a thorough survey about different definitions and patterns investigated in previous work. A prominent line of work [Shumailov et al., 2024, Kazdan et al., 2024, Bertrand et al., 2024, Suresh et al., 2024] focuses on the phenomenon of variance collapse, both empirically and theoretically demonstrating that models progressively lose information in the distribution’s tail, with variance tending toward zero. Another series of studies [Dohmatob et al., 2024a, Gerstgrasser et al., 2024, Bertrand et al., 2024, Dohmatob et al., 2024c, Fu et al., 2024] investigates model collapse through the lens of population risk and distributional shift, observing that the generated distribution increasingly diverges from the true data distribution, leading to rising population risk across recursive training cycles. Moreover, several works [Alemohammad et al., 2024a, Bohacek and Farid, 2023, Zhang et al., 2025b] report that models begin to generate hallucinated or unrealistic data as collapse progresses. Dohmatob et al. [2024d] further suggest that the inclusion of synthetic data alters the scaling laws of model performance. In addition, Cui et al. [2025] study the progressive mode collapse [Goodfellow et al., 2014] in diffusion models, where the number of modes in the generated distribution gradually decreases. In this paper, we introduce a novel perspective for analyzing model collapse in diffusion models by uncovering a generalization-to-memorization transition. We show that this transition is closely tied to the entropy of the training dataset, which serves as a key indicator of model generalizability. Our findings further motivate the development of entropy-based data selection strategies that effectively mitigate model collapse.

### v3

@@ 3.3  Finding III: The Correlation between Entropy and Generalization Score
We verify that the generalization score of the trained model is strongly correlated with the entropy of the training dataset. The similar collapsing trends of entropy and the generalization score motivate a deeper investigation into their relationship. In Figure˜4(a), we present a scatter plot of entropy versus generalization score across different dataset sizes and successive iterations, with the y-axis shown on a logarithmic scale. Notably, the entropy of the training dataset exhibits a significant linear relationship with the logarithm of the generalization score. The Pearson correlation coefficient is $0.91$ with a $p$-value near zero, quantitatively confirming the strength of this correlation. Furthermore, the scatter points corresponding to different dataset sizes are all approximately aligned along a single line, suggesting the generality of the relationship between entropy and generalization. Specifically, training datasets with higher entropy consistently yield better generalization in the trained model. We also observe that larger datasets typically have higher entropy and result in better generalization, aligning with the conclusion in [pmlr-v235-zhang24cn] regarding the connection between dataset size and generalization. For comparison, Figure˜4(b) shows the scatter plot of variance versus generalization score. The correlation appears substantially weaker than in Figure˜4(a), suggesting that variance in the training dataset may not directly influence the generalization performance of the trained model.
Conclusion.
@@ 4  Mitigating Model Collapse with Data Selection Methods
[…]
@@
[…]
@@ 5  Experiments
We empirically evaluate how our data selection strategies introduced in Section˜4 interact with the two self-consuming paradigms: replace and accumulate-subsample. In both settings, our method serves as a plug-in component for selecting high-quality and diverse training samples from the candidate pool. We demonstrate that the proposed strategies effectively alleviate memorization and reduce FID scores, thereby mitigating model collapse. Additionally, experiments on classifier-free guidance (CFG) [DBLP:journals/corr/abs-2207-12598] generation show that our method effectively mitigates diversity collapse of CFG.
@@
We conduct experiments on three widely used image generation benchmarks. CIFAR-10 [krizhevsky2009learning] consists of $32\times 32$ color images in $10$ classes. Due to computational constraints, we use a subset of 32,76832{,}768 training images. Our goal is not to achieve state-of-the-art FID among large diffusion models but rather to demonstrate that our method mitigates memorization in the self-consuming loop. As shown in Section˜3, this subset is sufficient to observe the transition from generalization to memorization. We also conduct experiments on subsets of FFHQ [DBLP:conf/pakdd/JainSH04], downsampled to $32\times 32$ resolution, and MNIST [DBLP:journals/pieee/LeCunBBH98], using 8,1928{,}192 and 12,00012{,}000 training images, respectively.
@@
For CIFAR-10 and FFHQ, we employ a UNet-based backbone [DBLP:conf/miccai/RonnebergerFB15] designed for $32\times 32$ RGB images, which predicts noise residuals. The network contains approximately 16M parameters. For MNIST, we use a similar UNet-based architecture adapted for single-channel inputs, with a total of 19M parameters. Detailed network configurations are provided in the Appendix.
@@
For iterative training and sampling, our implementation is based on the Hugging Face Diffusers codebase [von-platen-etal-2022-diffusers] of DDPM. We use a mixed-precision training of FP16 to train the models. We adopt an Adam optimizer with a learning rate of $10^{-4}$ and a weight decay of $10^{-6}$. The batch size is $128$. A $1000$-step denoising process is used, with all other hyperparameters set to their default values. For Threshold Decay Filter, we use an initial threshold of $60$ and a decay rate of $0.95$. We show in the Appendix that our method is robust in a wide range of hyperparameters.
@@
We use the generalization score and entropy to evaluate the effectiveness of our method in mitigating memorization. Additionally, we adopt the Fréchet Inception Distance (FID) [DBLP:conf/nips/HeuselRUNH17] as a metric to quantify the distributional divergence between generated images and real images.
@@ Results.
Diversity Improvement on Classifier-Free Diffusion Guidance (CFG) [DBLP:journals/corr/abs-2207-12598].
We further validate the effectiveness of our methods in the CFG setting and show that they substantially improve the diversity of generated images. CFG is a widely used conditional generation technique that consistently enhances perceptual quality but often sacrifices diversity [DBLP:conf/nips/ChidambaramGCL024, DBLP:journals/corr/abs-2207-12598, DBLP:journals/corr/abs-2505-19210]. Prior work by [yoon2024model] identifies the CFG scale as a key factor influencing the rate of model collapse and suggests that setting a moderate scale can help mitigate model collapse. Experimentally, we observe that the CFG generation can indeed generate clearer images than the unconditional baseline, as compared in Figures˜9(a) and 9(b). However, the diversity of generated images collapses rapidly even with a modest guidance scale, with samples within each class soon becoming nearly identical. In contrast, Figures˜9(c) and 9(d) demonstrate that augmenting CFG with our data selection methods significantly improves image diversity and yields substantially lower FID scores compared to the vanilla CFG paradigm. These results demonstrate that our approach effectively mitigates the diversity loss of CFG while preserving its quality advantage throughout the self-consuming loop.
@@
Figure 9: Comparisons of unconditional, CFG, and CFG augmented with filter method. (a)-(c): Generated samples on MNIST [DBLP:journals/pieee/LeCunBBH98] at the $8$-th iteration under the accumulate paradigm. The FIDs of images in (a)-(c) are $74.4$, $66.2$, and $\bm{22.4}$ respectively. (d): The estimated entropy of the generated dataset over iterations, which reflects the diversity of the generated images.
@@ 6  Conclusion and Discussion
Based on this work, we believe many future directions could be further investigated. This paper doesn’t rigorously explain why the entropy is collapsing. The finite dataset size, training and sampling errors, and the model bias could all contribute to the collapsing entropy (or the information loss). We envision that a more formal theoretical analysis of collapse dynamics could be developed based on theoretical models such as the mixture of Gaussians [DBLP:journals/corr/abs-2502-05743, DBLP:journals/corr/abs-2409-02426, DBLP:journals/corr/abs-2409-02374]. Besides, this framework could be extended to the language modality, investigating the discrete diffusion model [DBLP:conf/icml/LouME24, DBLP:journals/corr/abs-2506-07903, DBLP:journals/corr/abs-2502-06768]. Additionally, a more efficient algorithm is needed, as the current greedy selection method is computationally expensive.
Acknowledgement
LS, MW, HZ, ZZ, and QQ acknowledge funding support from NSF CCF-2212066, NSF CCF-2212326, NSF IIS 2402950, and ONR N000142512339, and the Google Research Scholar Award. MT is grateful for partial supports by NSF Grants DMS-1847802, DMS-2513699, DOE Grants NA0004261, SC0026274, and Richard Duke Fellowship. We also thank Prof. Peng Wang (University of Michigan and University of Macau), Xiang Li (University of Michigan), and Xiao Li (University of Michigan) for fruitful discussions and valuable feedback.
@@ A.1  Model Collapse
As state-of-the-art generative models continue to improve in image quality, AI-generated images have become increasingly indistinguishable from real ones and are inevitably incorporated into the training datasets of next-generation models. In fact, [DBLP:conf/iclr/AlemohammadCLHB24] have shown that the LAION-5B dataset [DBLP:conf/nips/SchuhmannBVGWCC22], which is used to train Stable Diffusion, indeed contains synthetic data. Unfortunately, recent studies [DBLP:journals/corr/abs-2408-16333, dohmatob2024strong, DBLP:journals/corr/abs-2410-16713, DBLP:conf/icml/DohmatobFYCK24, DBLP:conf/nips/DohmatobFK24, DBLP:conf/iclr/AlemohammadCLHB24, feng2024beyond] demonstrate that model performance deteriorates under such recursive training, potentially leading to model collapse, where the model generates increasingly homogeneous and meaningless content. The concept of model collapse was first introduced in [DBLP:journals/nature/ShumailovSZPAG24], which also provides a theoretical framework based on Gaussian models. Their analysis shows that if a Gaussian model is recursively estimated using data generated by its predecessor, its variance converges to zero—collapsing the distribution into a delta function concentrated at a single point.
@@
Following this important line of work, substantial research has further explored the model collapse phenomenon. [DBLP:journals/corr/abs-2412-17646] extend the theoretical analysis from Gaussian to Bernoulli and Poisson models. [DBLP:conf/nips/DohmatobFK24] study recursive training under high-dimensional linear and ridge regression settings, providing a linear error rate and proposing optimal regularization to mitigate collapse. [DBLP:conf/icml/DohmatobFYCK24] argue that the conventional scaling laws for foundation models break down when synthetic data is incorporated into the training set. [DBLP:conf/fat/WyllieSP24] demonstrate that model biases are amplified through recursive generation and proposes algorithmic reparation techniques to eliminate such biases and negative semantic shifts. [DBLP:journals/corr/abs-2404-05090] introduce a statistical model that characterizes the collapse process in language models and theoretically estimates the maximum allowable proportion of synthetic data before collapse occurs, validated empirically. [DBLP:conf/nips/FerbachBBG24] show that human preferences can be amplified in an iterative loop: modeling human curation as a reward process, the curated distribution $p_{t}$ converges to $p^{*}$ that maximizes the expected reward as $t\to\infty$. In the work of [DBLP:conf/nips/FerbachBBG24], the reward $r(x)$ is a pointwise function over individual images, whereas our method can be viewed as a curation strategy that maximizes dataset-level entropy at each iteration. Whether the theory in [DBLP:conf/nips/FerbachBBG24] extends to dataset-wise rewards $r(S)$, such as dataset entropy, remains an open question. Nonetheless, their intuition aligns with our results that entropy can be iteratively enhanced compared to vanilla methods. [DBLP:conf/iclr/FuWC0T25] provide the first generalization error bounds for Self-Consuming Training Loops (STLs), showing that mitigating model collapse requires maintaining a non-negligible portion of real data and ensuring recursive stability. In contrast, our work focuses on a specific collapse phenomenon—the generalization-to-memorization transition—and introduces an entropy-based data selection algorithm to mitigate this behavior. [zhu2024analyzing] provide a rigorous theoretical analysis for the model collapse in rectified flow models and then propose methods to mitigate it. Finally, [DBLP:journals/corr/abs-2501-03937] investigate collapse in diffusion models by analyzing the training dynamics of a two-layer autoencoder optimized via stochastic gradient descent, showing how network architecture shapes the model collapse.
@@
To mitigate model collapse, several strategies have been proposed in prior work. One common approach is to accumulate all previously generated samples along with real data into the training set—referred to as the accumulate paradigm in our paper. This strategy has been both empirically and theoretically validated. For example, [DBLP:journals/corr/abs-2404-01413] show that the accumulate paradigm prevents divergence of test error under a linear regression setup. [DBLP:journals/corr/abs-2410-22812] further establish the universality of the error upper bound across a broad class of canonical statistical models. In a related setting, [DBLP:journals/corr/abs-2410-16713] study an accumulate-subsample variant, confirms that the test error plateaus and examines interactions between real and synthetic data. [DBLP:conf/iclr/BertrandBDJG24] provide theoretical guarantees for stability in iterative training, assuming the initial model trained on real data is sufficiently accurate and the clean data proportion in each iteration remains high. Despite these findings, recent work [DBLP:journals/corr/abs-2410-04840] present a robust negative result, showing that model collapse persists even when real and synthetic data are mixed. Similarly, [DBLP:conf/iclr/AlemohammadCLHB24] argue that the accumulate paradigm merely delays, rather than prevents, collapse. Introducing fresh real data in each iteration may be necessary for long-term robustness. Other complementary approaches include verification [feng2024beyond] and re-editing [DBLP:journals/corr/abs-2412-14689]. For instance, [feng2024beyond] theoretically underscore the importance of data selection, though it does not propose a specific method. [DBLP:journals/corr/abs-2412-14689] target language models and introduces a token-level editing mechanism with theoretical guarantees. Compared to prior work, this paper proposes a novel entropy-based data selection method for diffusion models that improves both generalizability and image quality, thereby mitigating model collapse.
@@
Over the past few years, extensive research has explored model collapse from various perspectives and dimensions. An important study [DBLP:journals/corr/abs-2503-03150] makes a thorough survey about different definitions and patterns investigated in previous work. A prominent line of work [DBLP:journals/nature/ShumailovSZPAG24, DBLP:journals/corr/abs-2410-16713, DBLP:conf/iclr/BertrandBDJG24, DBLP:journals/corr/abs-2412-17646] focuses on the phenomenon of variance collapse, both empirically and theoretically demonstrating that models progressively lose information in the distribution’s tail, with variance tending toward zero. Another series of studies [DBLP:conf/nips/DohmatobFK24, DBLP:journals/corr/abs-2404-01413, DBLP:conf/iclr/BertrandBDJG24, DBLP:journals/corr/abs-2410-04840, DBLP:conf/icml/Fu [^0] T24] investigates model collapse through the lens of population risk and distributional shift, observing that the generated distribution increasingly diverges from the true data distribution, leading to rising population risk across recursive training cycles. Moreover, several works [DBLP:conf/iclr/AlemohammadCLHB24, DBLP:journals/corr/abs-2311-12202, DBLP:journals/corr/abs-2502-09963] report that models begin to generate hallucinated or unrealistic data as collapse progresses. [DBLP:conf/icml/DohmatobFYCK24] further suggest that the inclusion of synthetic data alters the scaling laws of model performance. In addition, [DBLP:journals/corr/abs-2501-03937] study the progressive mode collapse [DBLP:conf/nips/GoodfellowPMXWOCB14] in diffusion models, where the number of modes in the generated distribution gradually decreases. In this paper, we introduce a novel perspective for analyzing model collapse in diffusion models by uncovering a generalization-to-memorization transition. We show that this transition is closely tied to the entropy of the training dataset, which serves as a key indicator of model generalizability. Our findings further motivate the development of entropy-based data selection strategies that effectively mitigate model collapse.